OpenStudy (anonymous):

Al saves pennies. he agreed to give 3/7 of his pennies to bev if she would give 3/7 of what she got from al to carl. and if carl would give 3/7 of what he got from bev to dani. carl, and dani agreed and dani received 351 pennies. how many pennies did al have initially?

OpenStudy (anonymous):

I believe he had 4,459 pennies. I worked backwards and did 351/ (3/7) three times because that was how many time the pennies got shared among the people. And to check that answer just start off with 4459 and multiply it as many times it was shared.

OpenStudy (anonymous):

yay! thats what I got. thanks. i did the b=4/11xA and c=4/11xb